A man decides to build a road of 150 km in 300 days an employee 45 men. After 100 days,only 25 km of the road is completed,find how many extra men he needs to employ to finish the work in time
Answer by richwmiller(17219)

45*100=4500 man days
150/25=6 sections
one section is completed
You will need 5*4500 =22500 man days
22500/200 =112.5
112.5 men needed or 113 complete men
113-45=68 more men